OpenStudy (anonymous):

Evaluate the limit and justify each step by indicating the appropriate limit laws lim as u approaches -2 square root of (u^4 +3u +6)

OpenStudy (anonymous):

\[\lim_{u \rightarrow -2} \sqrt{u ^{4}+ 3u+6} \]

jimthompson5910 (jim_thompson5910):

replace u with -2 and evaluate like normal

OpenStudy (anonymous):

so is it 4

OpenStudy (anonymous):

and what limit laws were used

jimthompson5910 (jim_thompson5910):

yes the answer is 4 the limit law used was \[\Large \lim_{u\to c}(f(u)) = f(c) \] f(u) in this case is \[\Large f(u) = \sqrt{u^4+3u+6}\]

OpenStudy (anonymous):

does that limit law mean that the limit of a constant is the constant of its limit?

OpenStudy (anonymous):

and can u help me w/ anotehr limit problem

jimthompson5910 (jim_thompson5910):

no, that law is \[\Large \lim_{u\to c}(k) = k\]

OpenStudy (anonymous):

oh so what does that one taht u wrote mean verbally

OpenStudy (anonymous):

limu→c(f(u))=f(c)

jimthompson5910 (jim_thompson5910):

It's known as the direct substitution property

jimthompson5910 (jim_thompson5910):

your book/teacher may have another name for it

OpenStudy (anonymous):

oh okay...can u help me with anotehr limit problem

jimthompson5910 (jim_thompson5910):

sure

OpenStudy (anonymous):

so in this limit i evaluated it and i got 0 over 0 which is indeterminate. my teacher always says to somehow simplify teh problem if it is indeterminate \[\lim_{x \rightarrow -4} \sqrt{x ^{2}+9} - 5 / (x+4)\]

jimthompson5910 (jim_thompson5910):

are you familiar with L'Hospital ?

OpenStudy (anonymous):

no..what is taht

jimthompson5910 (jim_thompson5910):

ok nvm that option then

jimthompson5910 (jim_thompson5910):

what we need to do here is rationalize the numerator, how do we go about doing that?

OpenStudy (anonymous):

arent we supposed to multiply the top and bottom by teh conjugate (x-4)

jimthompson5910 (jim_thompson5910):

that's if we wanted to rationalize the denominator

jimthompson5910 (jim_thompson5910):

instead, we multiply top and bottom by \[\Large \sqrt{x^2+9}+5\]

jimthompson5910 (jim_thompson5910):

oh wait, made a typo

jimthompson5910 (jim_thompson5910):

sry, forgot something else doing that will give us \[\Large \frac{x^2-16}{(x+4)(\sqrt{x^2+9}+5)}\] \[\Large \frac{(x+4)(x-4)}{(x+4)(\sqrt{x^2+9}+5)}\] \[\Large \frac{x-4}{\sqrt{x^2+9}+5}\] from here, we can use the direct substitution property

jimthompson5910 (jim_thompson5910):

should be -4/5

jimthompson5910 (jim_thompson5910):

sry, made a typo earlier, but the work shown above is correct now

OpenStudy (anonymous):

oh yeah srry i made a mistake and taht is alright =)